PFI-ZEKE SPECTROSCOPY OF METAL-POLYCYCLIC AROMATIC HYDROCARBONS: Sc-PYRENE and Sc-NAPHTHALENE

\underline{JASON F. FULLER}, BRAD R. SOHNLEIN, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Gas-phase complexes of scandium with pyrene (C_{16}H_{20}) and naphthalene (C_{10}H_{8}) are produced by using laser ablation and a polycyclic aromatic hydrocarbon (PAH) oven in a pulsed molecular beam source. The resulting Sc-PAH complexes are characterized with pulsed field ionization-zero electron kinetic energy (PFI-ZEKE) photoelectron spectroscopy. The PFI-ZEKE spectrum of Sc-Pyrene exhibits vibrational progressions of 361 and 154 cm^{−1}and a vibrational spacing of 144 cm^{−1}, with the band origin at 40937 cm^{−1}. Through comparison with ab initio and Franck-Condon factor calculations, it is determined that this spectral profile arises from a Sc-pyrene complex in which the scandium atom is bound to the “outer” ring in an η^{4}- configuration. The 361 cm^{−1}progression is assigned to the Sc^{+}-ring stretch and 154 and 144 cm^{−1}features are assigned to pyrene “flying” modes in the cation and neutral, respectively. For Sc-naphthalene, the spectrum shows four vibrational intervals of 389, 347, 298 and 112 cm^{−1}, with the adiabatic ionization energy of 41192 cm^{−1}. A preliminary analysis indicated the Sc atom is in an η^{2} bonding mode in this complex. However, a detailed analysis is still in progress.